On a road map with a scale of 1/4 inch per 10 miles, the highway from Waukee to Winterset is 1 3/8 inches long. How many miles long is this highway?
A. 44
B. 55
C. 65
D. 70
E. 90

Found 2 solutions by doukungfoo, shauna broeske:
Answer by doukungfoo(195)

Ok, first we are told that:
On a road map with a scale of 1/4 inch per 10 miles
So this means that when measuring distance on the map every 1/4 ince is equal to 10 miles in the real world.
Next we are told:
the highway from Waukee to Winterset is 1 3/8 inches long
And we are asked:
How many miles long is this highway?


Ok so the first thing we have to do is figure out how my 1/4 inch sections are in 1 3/8 inches.
Well before I do anything towards solving the problem I am going to go ahead and change these fractions to decimal form. 1/4 = .25
and 1 3/8 = 1.375
Ok now how many times will .25 divide into 1.375
1.375/.25 = 5.5
So there are 5.5 one quarter inch sections in 1.375
Each one of these 5.5 is equal to 10 miles so.....
5.5 * 10 = 55
So the answer is 55 miles

Answer by shauna broeske(4)